# thanks: (No relevant financial or non-financial competing interests. Derived from preprint [23].)

### v3

@@ 2.1 The Extended Robust UCB Policy
Consier $K$ arms offering random rewards $\{X_{1},X_{2},\dots,X_{K}\}$ with distributions =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$. In contrast with the light-tailed class, the heavy-tailed distribution may yield a very high (or low) reward realization with a high probability. Consequently, the empirical mean estimator $\overline{X}_{i,s}=\frac{1}{s}\sum_{j=1}^{s}X_{i,j}$ for a sequence of i.i.d. random variables $\{X_{i,j}\}_{j=1}^{s}$ may not work well as an estimation to the expectation of $X_{i}$ for arm $i$. The appendix of Bubeck et al. [5] has shown that under the assumption of heavy-tailed distributions, with probability at least $1-\delta$,
@@
where $1<r\leq 2$. This bound is too loose to obtain a UCB policy achieving the logarithmic regret growth. To remedy this issue of divergence, we adopt the median-of-means estimator proposed in Bubeck et al. [5].
@@
Given an arbitrary $\varepsilon>0$. Suppose that
$\log\delta^{-1}>1/\varepsilon,\ s\geq(8+\varepsilon)\log\delta^{-1}\frac{8\log(\delta^{-1})+1}{\varepsilon\log(\delta^{-1})-1}$
and $v=\mathbb{E}[|X_{i}-\theta_{i}|^{p}]<\infty$, for some $p$ such that $p>1$. Then with probability at least $1-\delta$, the median-of-means estimator
$\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},\lceil 8\log\delta^{-1}\rceil)\geq\theta_{i}-(12v)^{1/p}\left(\frac{(8+\varepsilon)\log(\delta^{-1})}{s}\right)^{(p-1)/p}.$
Also, with a probability of at least $1-\delta$,
$\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},\lceil 8\log\delta^{-1}\rceil)\leq\theta_{i}+(12v)^{1/p}\left(\frac{(8+\varepsilon)\log(\delta^{-1})}{s}\right)^{(p-1)/p}.$
Proof.
$k$ $:=\lceil 8\log\delta^{-1}\rceil,$ (8)
$\chi$ $:=\frac{(8+\varepsilon)\log\delta^{-1}}{s},$ (9)
$\eta$ $:=(12v)^{1/p}\chi^{(p-1)/p},$ (10)
According to the appendix of Bubeck et al. [5],
Note that $N=\lfloor s/\lceil 8\log\delta^{-1}\rceil\rfloor$. Since $s\geq(8+\varepsilon)\log\delta^{-1}\frac{8\log(\delta^{-1})+1}{\varepsilon\log(\delta^{-1})-1}$, a direct computation yields $\xi\leq 1/4$. Then using Hoeffding’s inequality for the tail of a binomial distribution [14], we have
Similarly, the other inequality also holds by symmetry. ∎
With the help of the median-of-means estimator, we can now define the upper confidence bound for the extended robust UCB policy. As mentioned above, we assume that for any arm distribution $X_{i}\in\mathcal{F}$ with mean $\theta_{i}$, there is a moment control coefficient $C_{p,q}$ ($1<q<p$) such that
@@
Recall the notations introduced by (8) and (9). Given an arbitrary $\varepsilon>0$, a moment order $p>1$, another moment order $1<q<p$, and the moment control coefficient $C_{p,q}$, define the upper confidence bound of the extended robust UCB as (value of the fraction is set to +$\infty$+ ∞ if divided by 00)
@@
Proof.
@@
The proof is thus completed by the definition of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)$. ∎
Corollary 1.
$\hat{\mu}(\{Y_{i,j}\}_{j=1}^{s},k)-v_{q}\leq C^{\prime}v_{q}\chi^{(p-q)/p}.$
Proof.
This corollary is a direct consequence of Lemma 1 and the inequality in (16). ∎
Finally, with the upper confidence bound estimator $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s})$, we complete the extended robust UCB policy as follows.
1 Input: $\varepsilon>0$, $K$: the number of arms
2 Initialize: $s_{i}\leftarrow 0$ and $\{X_{i,j}\}_{j=1}^{s_{i}}$ an empty sequence for each arm $i$, $1\leq i\leq K$
3 for $t\leftarrow 1$ to $T$ do
4       for $i\leftarrow 1$ to $K$ do
5             if $s_{i}=0$ then
6                  Set arm $i$’s upper confidence bound to +$\infty$+ ∞
7             end if
8            else
9                  Compute arm $i$’s upper confidence bound as $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$
10             end if
11
12       end for
13      Choose an arm $j$ that maximizes the upper confidence bound and obtain reward $x$
14       Update: $s_{j}\leftarrow s_{j}+1$ and append $x$ to sequence $\left\{X_{j,k}\right\}_{k=1}^{s_{j}}$
15 end for
Algorithm 1 The Extended Robust UCB Policy
Theorem 1.
Suppose $p>1$ such that the $p$-th order moments exist for all $F_{i}\in\mathcal{F}$, $1<q<p$ and $C_{p,q}$ such that (12) holds. Then for any $\varepsilon>0$, the regret $\mathcal{R}_{\pi^{*}}^{\mathcal{F}}(T)$ of the extended robust UCB policy $\pi^{*}$ has the logarithmic order with respect to the time horizon $T$.
To prove Theorem 1, we establish the following core lemma to bound the time $s_{i}$ spent on each arm $i$ by the extended robust UCB policy.
Lemma 3.
The expected number of times $\mathbb{E}[s_{i}(T)]$ that a sub-optimal arm is chosen by the extended robust UCB has an upper bound of the logarithmic order with $T$.
Proof.

### v4

@@ 2.1 The Extended Robust UCB Policy
Consider $K$ arms offering random rewards $\{X_{1},X_{2},\dots,X_{K}\}$ with heavy-tailed distributions =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$. In contrast with the light-tailed class, a heavy-tailed distribution may yield a very high (or low) reward realization with a high probability. Consequently, the empirical mean estimator $\overline{X}_{i,s}=\frac{1}{s}\sum_{j=1}^{s}X_{i,j}$ for a sequence of i.i.d. random variables $\{X_{i,j}\}_{j=1}^{s}$ may not work properly to estimate the expectation of $X_{i}$ for arm $i$. The appendix of  Bubeck et al. [5] has shown that under the assumption of heavy-tailed distributions, with probability at least $1-\delta$,
@@
where $1<r\leq 2$. This bound is too loose to obtain a UCB policy for achieving logarithmic regret growth. To remedy this issue, we adopt the median-of-means estimator proposed in  Bubeck et al. [5].
@@
Given any $\epsilon>0$. Suppose that
$\log\delta^{-1}>1/\epsilon,\ s\geq(8+\epsilon)\log\delta^{-1}\frac{8\log(\delta^{-1})+1}{\epsilon\log(\delta^{-1})-1}$
and $v=\mathbb{E}[|X_{i}-\theta_{i}|^{p}]<\infty$ for some $p$ such that $p>1$. Then with probability at least $1-\delta$, the median-of-means estimator
$\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},\lceil 8\log\delta^{-1}\rceil)\geq\theta_{i}-(12v)^{1/p}\left(\frac{(8+\epsilon)\log(\delta^{-1})}{s}\right)^{(p-1)/p}.$
Also, with probability at least $1-\delta$,
$\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},\lceil 8\log\delta^{-1}\rceil)\leq\theta_{i}+(12v)^{1/p}\left(\frac{(8+\epsilon)\log(\delta^{-1})}{s}\right)^{(p-1)/p}.$
With the help of the median-of-means estimator, we can now define the upper confidence bound for the extended robust UCB policy. As mentioned in (5), we first assume that for any arm distribution $X_{i}\in\mathcal{F}$ with mean $\theta_{i}$, there is a known moment control coefficient $C_{p,q}$ ($1<q<p$) such that
@@
Let $k=\lceil 8\log\delta^{-1}\rceil,\chi=\frac{(8+\epsilon)\log\delta^{-1}}{s}$. Given an arbitrary $\epsilon>0$, a moment order $p>1$, another moment order $1<q<p$, and the moment control coefficient $C_{p,q}$, define the upper confidence bound of the extended robust UCB as (value of the fraction is set to +$\infty$+ ∞ if divided by 00)
@@
Corollary 1.
$\hat{\mu}(\{Y_{i,j}\}_{j=1}^{s},k)-v_{q}\leq C^{\prime}v_{q}\chi^{(p-q)/p}.$
Please see (27) for the definitions of $Y_{i,j},v_{q},Y_{i},v_{p}$. Finally, with the upper confidence bound $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s})$, we complete the extended robust UCB policy as follows.
Input: $\epsilon>0$, $K$: the number of arms
Initialize: $s_{i}\leftarrow 0$ and $\{X_{i,j}\}_{j=1}^{s_{i}}$ an empty sequence for each arm $i$, $1\leq i\leq K$
for $t\leftarrow 1$ to $T$ do
       for $i\leftarrow 1$ to $K$ do
             if $s_{i}=0$ then
                  Set arm $i$’s upper confidence bound to +$\infty$+ ∞
             end if
            else
                  Compute arm $i$’s upper confidence bound as $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$
             end if
       end for
      Choose an arm $j$ that maximizes the upper confidence bound and obtain reward $x$
       Update: $s_{j}\leftarrow s_{j}+1$ and append $x$ to sequence $\left\{X_{j,k}\right\}_{k=1}^{s_{j}}$
end for
Algorithm 1 The Extended Robust UCB Policy $\pi^{*}$
Theorem 1.
Suppose $p>q>1$ such that the $p$-th order moments exist for all $F_{i}\in\mathcal{F}$ and (6) is satisfied with known $C_{p,q}$. Then for any $\epsilon>0$, the regret $\mathcal{R}_{\pi^{*}}^{\mathcal{F}}(T)$ of the extended robust UCB policy $\pi^{*}$ has the logarithmic order with respect to the time horizon $T$.
To prove Theorem 1, we just need to show the following core lemma to bound the time $s_{i}$ spent on each suboptimal arm $i$ by the extended robust UCB policy.
Lemma 3.
The expected number of times $\mathbb{E}[s_{i}(T)]$ that a suboptimal arm is chosen by the extended robust UCB policy $\pi^{*}$ has an upper bound of the logarithmic order with $T$.
A Monte Carlo simulation example based on the Pareto distribution of type III (see Sec. 1.1) is shown in Fig. 1. Note that the regret rate grows linearly at the beginning to prepare the ground for ranking until at least one arm has a finite UCB (i.e. $\chi$ defined in (24) becomes sufficiently small such that, within the $\sup$ operator of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$, both the denominator is not zero and the growth rate of the right-hand side with $\theta$ becomes small). Note that if $C_{p,q}$ is unknown, then the player can simply choose any increasing function $f$ with $f(t)\rightarrow+\infty$ as $t\rightarrow+\infty$ to replace $C_{p,q}$ and achieve a regret arbitrarily close to logarithmic order without requiring any knowledge of the reward moments (see Appendix F for a proof).
Figure 1: The performance of extended robust UCB
2.2 Further Analysis and Comparison of The Extended Robust UCB
Now we consider scenarios of specific chosen $p$ and $q$ to further improve the efficiency of the extended robust UCB policy.
Case 1. Suppose that the $p^{\prime}$-th order of moments of reward distributions exists with $p^{\prime}>2$ for all arms. Then we can choose $p$ and $q$ such that $p=2q\leq p^{\prime}$. The inequality (29) can be refined to
$\begin{split}\mathbb{E}[|Y_{i}-v_{q}|^{p/q}]=\mathbb{E}Y_{i}^{2}+v_{q}^{2}-2v_{q}\mathbb{E}Y_{i}=\mathbb{E}Y_{i}^{2}-v_{q}^{2}=v_{p}-v_{q}^{2}\leq(C_{p,q}-1)v_{q}^{p/q}.\end{split}$ (8)
Hence, we can refine $C^{\prime}$ from (7) to
$C^{\prime}=\sqrt{12(C_{p,q}-1)},$ (9)
which can be much smaller than the original definition (7).
Now we look at the definition of $\chi_{M}$ in Appendix E that ensures the three inequalities (39), (40) and (41) to hold. The refinement of $C^{\prime}$ leads to a smaller right-hand side for all these inequalities, i.e. the constant $\chi_{M}$ will be larger and the regret upper bound will be smaller, which further implies that the policy’s selection will converge to the best arm faster.
Case 2. Assume that the $4$-th order of moments of reward distributions exists as in  Lattimore [22]. Then we can choose $p=4$ and $q=2$. In this case, our UCB in Definition 2 can be rewritten as
[…]
In other words, $C_{p,q}$ degenerates to an upper bound on kurtosis. The function $B(x)$ defined in (36) becomes
and the three inequalities (39), (40) and (41) are refined to
11 11 $>C^{\prime}\sqrt{x},$ (10)
11 11 $>48x\tau(x),$ (11)
$\Delta_{i}$ $\geq\sqrt{v_{2}x}\left(\sqrt{B(x)}+2\sqrt{3}\right).$ (12)
Under these refined estimations for $\chi_{M}$, the leading coefficient of $\log T$ in (46) is significantly decreased. Now we compare the extended robust UCB and the scale free algorithm proposed in  Lattimore [22], under the same assumption specified at the beginning of Case 2.
We first address a flaw of Lemma 2 in Bubeck et al. [5]. This lemma, crucial in Lattimore [22] for deriving the regret upper bound of the scale free algorithm, is examined more closely below. For clarity and coherence with the notation employed in this paper, we present the following adapted version.
A Flawed Statement  Bubeck et al. [5]: Let $\delta\in(0,1)$ and $p\in(1,2]$. Let $\{X_{i,j}\}_{j=1}^{s}$ be i.i.d. random variables with mean $\mathbb{E}X_{i}=\theta_{i}$ and (centered) $p$-th moment $v_{p}$. Let $k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor$. Then with probability at least $1-\delta$,
The last inequality in the proof of this statement in  Bubeck et al. [5] assumes that $\exp(-k^{\prime}/8)\leq\delta$, equivalent to $k^{\prime}\geq 8\log\delta^{-1}$. However, the definition of $k^{\prime}$ in this statement is
$k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor.$
In the following, we give the corrected version of the above statement.
Lemma 4.
Let $\delta\in(0,1)$ and $p\in(1,2]$. Let $\{X_{i,j}\}_{j=1}^{s}$ be i.i.d. random variables with mean $\mathbb{E}X_{i}=\theta_{i}$ and (centered) $p$-th moment $v_{p}$, where $s$ is chosen such that $\lfloor s/2\rfloor\geq 8\log\delta^{-1}$. Let $k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor$. Then with probability at least $1-\delta$,
The next theorem shows that with a properly chosen parameter $\epsilon>0$, the upper bound given in (46) is also tighter than that in Lattimore [22] in long run if $\Delta_{i}$ is small, i.e., arms are hard to be distinguished (the low-discrimination case). In the following theorem, we formally state this result while correcting the upper bound in Lattimore [22] by Lemma 4.
Theorem 2.
Choose any $8<\epsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$ ($\approx 44.158$) and let $\delta=t^{-2}$. The upper bound given in Lattimore [22] of $\mathbb{E}[s_{i}(T)]$ is greater than that given in (46) for sufficiently large $T$ and sufficiently small $\Delta_{i}$. That is, after dividing both regret upper bounds by $\log\delta^{-1}$ and letting $T\to+\infty$, we have
$(8+\epsilon)\left(\frac{8}{\epsilon}+\max\left\{\frac{8}{\epsilon},\frac{1}{\chi_{M}}\right\}\right)<3648\max\left\{\frac{(C^{\prime})^{2}}{12},\frac{v_{2}}{\Delta_{i}^{2}}\right\}+16$ (13)
for sufficiently small $\Delta_{i}$. The regret upper bound of the scale free algorithm proposed in  Lattimore [22] is thus asymptotically larger than that of the extended robust UCB policy for $\epsilon\in(8,\frac{280-16\sqrt{2}}{2\sqrt{2}+3})$.
Last, we show that the regret of the extended robust UCB deviates from the theoretical lower bound (in the non-parametric setting) by only a constant factor and a constant term as in the scale free algorithm of  Lattimore [22]. First, we restate the lower bound derived in  Lattimore [22] as follows.
Theorem 3.
Let $H_{\kappa_{0}}$ be the set of distributions that have kurtosis less than $\kappa_{0}$. Assume $\Delta>0$ and $\kappa_{0}\geq 7/2$. Suppose that $X\in H_{\kappa_{0}}$ has a mean of $\mu$, a positive variance of $\sigma^{2}$, and a kurtosis of $k$. Then
[…]
Based on this, we can draw the following conclusion:
Theorem 4.
Let $\delta=t^{-2}$. Under Case 2, the regret of the extended robust UCB policy applied to $H_{\kappa_{0}}$ ($\kappa_{0}\geq 7/2$) differs from the lower bound of regret by a constant factor and a constant term.
3 Light-Tailed Reward Distributions
We point out that our policy and results also work for the light-tailed distributions since our construction depends only on the existence of moments. However, the leading constant of the logarithmic regret order can be improved by our past work in Liu and Zhao [24], which we formally and briefly discuss here with a comparison with some subsequent work following Liu and Zhao [24].
3.1 The UCB1-LT Policy
The empirical mean is adopted in estimating the upper confidence bound for the class of light-tailed reward distributions. As before, let $\overline{X}_{i,s}$ denote the empirical mean $\frac{1}{s}\sum_{j=1}^{s}X_{i,j}$ for arm $i$, where $\{X_{i,j}\}_{j=1}^{s}$ form the i.i.d. random reward sequence drawn from an unknown distribution $X_{i}\in\mathcal{F}$.
Lemma 5.
(Bernstein-type bound) For i.i.d. random variables $\{X_{i,j}\}_{j=1}^{s}$ drawn from a light-tailed distribution $X_{i}$ (with mean $\theta_{i}$) with a finite moment-generating function $M(u)$ over range $u\in[-u_{0},u_{0}]$ for some $u_{0}>0$. We have, $\forall\epsilon>0$,
$\mathbb{P}(\overline{X}_{i,s}-\theta_{i}\geq\epsilon)\leq\begin{cases}\exp(-\frac{s}{2\zeta}\epsilon^{2}),&\epsilon<\zeta u_{0}\\ \exp(-\frac{su_{0}}{2}\epsilon),&\epsilon\geq\zeta u_{0}\\ \end{cases},$ (15)
where $\zeta>0$ satisfies $\zeta\geq\sup_{|u|\leq u_{0}}{M^{(2)}(u)}$. A similar bound also holds for $\mathbb{P}(\overline{X}_{i,s}-\theta_{i}\leq-\epsilon)$ by symmetry.
Input: $a_{1}\geq 8\zeta$, $a_{2}\geq a_{1}/(\zeta u_{0})$, $K$: the number of arms
Initialize: $s_{i}\leftarrow 0$ and $\overline{X}_{i}\leftarrow 0$ for each arm $i$, $1\leq i\leq K$
for $t\leftarrow 1$ to $T$ do
       for $i\leftarrow 1$ to $K$ do
             if $s_{i}=0$ then
                  Assign +$\infty$+ ∞ to the upper confidence bound
             end if
            else
                   if $\sqrt{\frac{a_{1}\log t}{s_{i}}}<\zeta u_{0}$ then
                         Compute the upper confidence bound as $\overline{X}_{i}+\sqrt{\frac{a_{1}\log t}{s_{i}}}$
                   end if
                  else
                         Compute the upper confidence bound as $\overline{X}_{i}+\frac{a_{2}\log t}{s_{i}}$
                   end if
             end if
       end for
      Choose an arm $j$ that maximizes the upper confidence bound and obtain reward $x$
       $\overline{X}_{j}\leftarrow\frac{x+s_{j}\overline{X}_{j}}{s_{j}+1}$ and $s_{j}\leftarrow s_{j}+1$
end for
Algorithm 2 The UCB1-LT Policy
The UCB1-LT policy considers two upper confidence bounds and alternatively uses one of them according to values of $t$ and $s_{i}(t-1)$. To minimize the theoretical regret upper bound, we can choose $a_{1}=8\zeta$ and $a_{2}=\frac{8}{u_{0}}$. The following theorem shows that UCB1-LT achieves logarithmic regret growth for the light-tailed reward distributions.
Theorem 5.
For the light-tailed reward distributions, the regret of the UCB1-LT policy satisfies the following inequality:
$\mathcal{R}_{\pi}^{\mathcal{F}}(T)\leq\sum_{i:\theta_{i}<\theta_{1}}\Delta_{i}\left(\max\left\{\frac{4a_{1}}{\Delta_{i}^{2}},\frac{2a_{2}}{\Delta_{i}}\right\}\log T+1+\frac{\pi^{2}}{3}\right).$ (16)
In Appendix A, we compare our UCB1-LT policy with the $(\alpha,\psi)$-UCB policy proposed in Bubeck and Cesa-Bianchi [4].
4 Conclusion and Acknowledgment
In this paper, we have proposed two order-optimal UCB policies, namely the extended robust UCB and UCB1-LT, dealing with the heavy-tailed and light-tailed reward distributions in the frequentist multi-armed bandit problems, respectively. We are grateful for the help from Mr. Haoran Chen during the initial stage of this project, and Dr. Yaoqing Yang from Dartmouth College in improving this article.
Appendix A From UCB1-LT to $(\alpha,\psi)$-UCB
After the establishment of UCB1-LT by Liu and Zhao [24],  Bubeck and Cesa-Bianchi [4] subsequently proposed the $(\alpha,\psi)$-UCB under a more general assumption on the moment generating functions. Specifically, the $(\alpha,\psi)$-UCB policy assumes that the distribution of reward $X_{i}\in\mathcal{F}$ satisfies the following condition: there exists a convex function $\psi$ defined on $\mathbb{R}$ such that for all $\lambda>0$,
For example, if $X_{i}$ is bounded in $[0,1]$ as assumed in UCB1, one can choose $\psi(\lambda)=\frac{\lambda^{2}}{8}$ and (17) becomes the well-known Hoeffding’s lemma [4]. Since +$\infty$+ ∞ is allowed in the range of $\psi$, this assumption is more general than that in UCB1-LT. The Legendre-Frenchel transform $\psi^{*}$ of $\psi$ is defined as
$\psi^{*}(\epsilon):=\sup_{\lambda\in\mathbb{R}}\lambda\epsilon-\psi(\lambda).$
Theorem 6 (Bubeck and Cesa-Bianchi [4]).
For the UCB1-LT policy, we can choose
$\psi^{*}(\epsilon)=\begin{cases}\frac{\epsilon^{2}}{2\zeta},&|\epsilon|<\zeta u_{0}\\ u_{0}(|\epsilon|-\frac{\zeta u_{0}}{2}),&|\epsilon|\geq\zeta u_{0}\end{cases}.$
Thus the inverse $(\psi^{*})^{-1}(x)$ of $\psi^{*}$ for $x\geq 0$ will be
If we choose $a_{1}=8\zeta$ and $a_{2}=\frac{8}{u_{0}}$ in UCB1-LT, we get the same UCB under the condition $\frac{\log t}{s_{i}(t-1)}<\frac{\zeta u_{0}^{2}}{8}$. For the case $\frac{\log t}{s_{i}(t-1)}\geq\frac{\zeta u_{0}^{2}}{8}$, the upper confidence bound of $(\alpha,\psi)$-UCB is not larger than that of UCB1-LT and thus may achieve a smaller theoretical upper bound of regret as follows:
where
Lemma 6.
Choose the parameter $\alpha$ in $(\alpha,\psi)$-UCB such that $\alpha\leq 4$, and assume that Assumption (4) is satisfied and thus $\psi(\lambda)$ is chosen as in (20). Then the logarithmic regret upper bound in Theorem 6 has a leading constant no larger than that in Theorem 5, i.e.
Appendix B Proof of Lemma 1
$k$ $:=\lceil 8\log\delta^{-1}\rceil,$ (23)
$\chi$ $:=\frac{(8+\epsilon)\log\delta^{-1}}{s},$ (24)
$\eta$ $:=(12v)^{1/p}\chi^{(p-1)/p},$ (25)
According to the appendix of  Bubeck et al. [5],
Note that $N=\lfloor s/\lceil 8\log\delta^{-1}\rceil\rfloor$. Since $s\geq(8+\epsilon)\log\delta^{-1}\frac{8\log(\delta^{-1})+1}{\epsilon\log(\delta^{-1})-1}$, a direct computation yields $\xi\leq 1/4$. Then using Hoeffding’s inequality for the tail of a binomial distribution Hoeffding [15], we have
Similarly, the other inequality also holds by symmetry. $\hfill\qed$
Appendix C Proof of Lemma 2
@@
The proof is thus completed by the definition of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)$. $\hfill\qed$
Appendix D Proof of Corollary  1
This corollary is a direct consequence of Lemma 1 and the inequality in (29). $\hfill\qed$
Appendix E Proof of Lemma 3